Help!
A chemist needs a solution of tannic acid that is 70% pure. How much distilled water must she add to 5 gallons of acid which is 90% pure to obtain 70% solution?

Answer by longjonsilver(2297)

look at my Lesson on "Mixture" questions.

BEFORE: 5 gallons 90% acid and x gallons of 0% acid
AFTER: (5+x) gallons 70% acid

5*0.90 + x*0.00 = (5+x)*0.70
4.5 + 0 = 3.5 + 0.7x
1 = 0.7x
x = 1/0.7
x = 1.43 gallons of distilled water (to 2dp)

jon.